# Deep GraphRAG: A Balanced Approach to Hierarchical Retrieval and Adaptive Integration

### v2

@@ 2.1.1. Graph Construction and Hierarchy
Entity Resolution: To maintain topological consistency, we implement a hybrid resolution strategy. Initial candidate pairs for merging are identified by computing cosine similarity on entity descriptions embedded via the bge-m3 model. Pairs exceeding a strict similarity threshold ($\tau>0.95$) undergo a secondary verification step where an LLM acts as a discriminator to confirm if the nodes refer to the same real-world concept (e.g., merging ”U.S.” and ”United States”).
@@ 2.1.2. Retrieval Process
The retrieval (Algorithm 1) employs a beam search ($k=3$) to traverse $\mathcal{C}$ from top to bottom. Relevance at each level is scored using cosine similarity between $q$ and $D(\cdot)$. Finally, retrieved knowledge is structured hierarchically (community summaries, then entity details) for LLM integration.
@@
Input: Query $q$, Community $C$, Beam width $k$, Top-$m$ entities $m$ Output: Final result set $R$
1:Phase 1: Top-level Search (Beam Search, $k=3$)
2: Score $C_{top}$ via $\text{sim}_{\cos}(q,D(c_{top}))$; $C_{mid}\leftarrow\text{Top-k}$
3:Phase 2: Middle-level Search (Beam Search, $k=3$)
4: Expand $C_{mid}$ to $C^{\prime}_{mid}$; Score via $\text{sim}_{\cos}(q,D(c^{\prime}_{mid}))$
[…]
6:Phase 3: Entity-level Search
7: Expand $C_{stop}$ to $V_{cand}$; Score via $\text{sim}_{\cos}(q,D(v))$
[…]
9:Phase 4: Knowledge Integration
[…]
11:return $R$
$2.$ Dynamic Weighting Reward GRPO (DW-GRPO)
@@ 2.2.1. DW-GRPO in Deep GraphRAG
    Relevance ($r_{\text{rel}}$): Measures how well $C$ answers $Q$. We use a pre-trained cross-encoder model $f_{\text{cross}}$ to score the query-output pair:
@@
    Faithfulness ($r_{\text{faith}}$): Measures the semantic fidelity of $C$ to the original knowledge $K$. We use the F1-score from BERTScore $f_{\text{BERT}}$:

### v3

@@ 2.1.1. Graph Construction and Hierarchy
Entity Resolution: To maintain topological consistency, we implement a hybrid resolution strategy. Initial candidate pairs for merging are identified by computing cosine similarity on entity descriptions embedded via the bge-m3(Chen et al., 2023) model. Pairs exceeding a strict similarity threshold ($\tau>0.95$) undergo a secondary verification step where an LLM acts as a discriminator to confirm if the nodes refer to the same real-world concept (e.g., merging ”U.S.” and ”United States”).
@@ 2.1.2. Retrieval Process
The retrieval (Algorithm 1) employs a coarse-to-fine beam search ($k=3$) to traverse $\mathcal{C}$ from top to bottom. To balance efficiency and accuracy, we adopt a hybrid scoring strategy: relevance at the top coarse level is quickly estimated using re-ranker, while the intermediate expansion utilizes a bge-reranker-v2-m3 to capture subtle semantic alignments between the query $q$ and community summaries.
At the final entity level, we ensure topological awareness by constructing context-aware representations. Specifically, the representation of a candidate entity $v$ is dynamically formed by concatenating its local embedding with its parent community’s vector (i.e., $D_{ctx}(v)=[D(v);D(c_{parent})]$) before final scoring.
@@
Input: Query $q$, Community $C$, Beam width $k$, Top-$m$ entities $m$
Output: Final result set $R$
1:Phase 1: Top-level Search (Coarse Filtering)
2:   Score $C_{top}$ via $\text{Re-ranker}(q,D(c_{top}))$
3:   $C_{mid}\leftarrow\text{Top-k results from }C_{top}$
4:Phase 2: Middle-level Search (Fine-grained Re-ranking)
5:   Expand $C_{mid}$ to sub-communities $C^{\prime}_{mid}$
6:   Score $C^{\prime}_{mid}$ via $\text{Re-ranker}(q,D(c_{mid})+D(c^{\prime}_{mid}))$
7:   $C_{stop},V_{cand}\leftarrow\text{Top-k results from }C^{\prime}_{mid}$
8:Phase 3: Entity-level Search (Context-Aware)
9:   Expand $C_{stop}$ to candidate entities $V_{cand}$
10:for each entity $v\in V_{cand}$ with parent $c_{parent}$ do
11:    $D_{ctx}(v)\leftarrow\text{Concat}(D(v),D(c_{parent}))$
12:    $Score(v)\leftarrow\text{sim}_{\cos}(q,D_{ctx}(v))$
13:end for
14:   $R_{pre}\leftarrow\text{Top-m results from }V_{cand}$ based on $Score(v)$
15:Phase 4: Knowledge Integration
16:   $R\leftarrow\text{Hierarchical-Integration}(R_{pre})$
17:return $R$
2.2. Dynamic Weighting Reward GRPO (DW-GRPO)
@@ 2.2.1. DW-GRPO in Deep GraphRAG
    Relevance ($r_{\text{rel}}$): Measures how well $C$ answers $Q$. We use a pre-trained cross-encoder model (bge-reranker-v2-m3) $f_{\text{cross}}$ to score the query-output pair:
@@
    Faithfulness ($r_{\text{faith}}$): Measures the semantic fidelity of $C$ to the original knowledge $K$. We use the F1-score from BERTScore (bge-m3) $f_{\text{BERT}}$: